OpenStudy (anonymous):

Solve the following proportion. (Dieing of Boredom .-.) r/5 = 9/21 a. r = 9 b. r = 2.14 c. r = 11.66 d. r = 37.8

OpenStudy (solomonzelman):

\(\LARGE\color{blue}{ \frac{r}{5}=\frac{9}{21} }\) cross multiply, \(\Large\color{blue}{ r \times 21=9 \times 5 }\)

OpenStudy (solomonzelman):

Simplify this expression and solve for x. Tell me if you need more help.....

OpenStudy (solomonzelman):

you are correct.

OpenStudy (solomonzelman):

2.14 is the correct answer....

OpenStudy (anonymous):

In 10 months, 62.6 gallons were used. Find the unit rate. a. 0.16 gallons/month b. 10 gallons/month c. 6.3 gallons/month d. 626 gallons/month

OpenStudy (anonymous):

i was gonna say d. if it had a decimal between 6 and 26

OpenStudy (solomonzelman):

you need the rate per month. So if in 10 months 62.6 gallons were used then, in 1 month you used 10 times less. So you divide 62.6 by 10. (The answer is not there, you will need to round it up to tenths) .

OpenStudy (anonymous):

Write the ratio as a fraction in simplest form. 8 out of 10 a. 8/10 b. 5/4 c. 2 d. 4/5